ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 95, 1996

Agranovsky, M. L. and A. M. Semenov Deformations of balls in Schiffer’s conjecture for Riemannian symmetric spaces

43

Bazzoni, S. and L. Salce Groups in the class semigroups of valuation domains

135

Bican, L. and K. M. Rangaswamy PB^{∞}-subgroups of Butler groups

375

Boyle, M. and D. Handelman Orbit equivalence, flow equivalence and ordered cohomology

169

Burton, R. M., M. Denker and M. Smorodinsky Finite state bilaterally deterministic strongly mixing processes

115

Casazza, P. G. and N. J. Kalton Unconditional bases and unconditional finite-dimensional decompositions in Banach spaces

349

Denker, M. See R. M. Burton

Dumont, J-M., T. Kamae and S. Takahashi Minimal cocycles with the scaling property and substitutions

393

Ginzburg, D. A Rankin–Selberg integral for the adjoint L-function

of Sp_{4} 301

Handelman, D. See M. Boyle

Kalton, N. J. See P. G. Casazza

Kamae, T. See J-M. Dumont

Mormul, P. and M. Zhitomirskii Modules of vector fields, differential forms and degenerations of differential systems

411

Morrisey, T. J. A degenerate Hartman theorem  157

Nakao, M. Decay of solutions of the wave equation with a local degenerate dissipation

25

Peres, Y. Points of increase for random walks  341

Rangaswamy, K. M. See L. Bican

Reichstein, Z. and N. Vonessen Rational central simple algebras  253

Salce, L. See S. Bazzoni

Sankey, A. D. Regular weights of full rank on strongly regular graphs  1

Semenov, A. M. See M. L. Agranovsky

Serre, J-P. Two letters on quaternions and modular forms (mod p) 281

Shelah, S. Further cardinal arithmetic  61

Silver, D. S. and S. G. Williams Augmented group systems and shifts of finite type

231

Sjögren, P. and F. Soria Weak type (1,1) estimates for some integral operators related to rough maximal functions

211

Smorodinsky, M. See R. M. Burton Soria, F. See P. Sjögren Takahashi, S. See J-M. Dumont Vonessen, N. See Z. Reichstein Williams, S. G. See D. S. Silver Zhitomirskii, M. See P. Mormul